A kennel owner has 164 ft of fencing to enclose a rectangular region. He wants to subdivide it into 3 sections of equal length. If the total area of the enclosed region is 576 square ft what are the dimensions.
I know that the answer is 18 ft by 32 ft or 64 ft by 9ft but not how to get it
Answer #1
Since there are three sections of equal length, there must be 4 lengths (2 outside and 2 inside) and the width on both sides.

4L + 2W = 164
2W = 164 - 4L
W = 82 -2L

L * W = 576

Substitute 82 - 2L for W in the second equation and solve for L. Put that value in the first equation and solve for W. Check by putting both values in the second equation.

I hope this helps. Thanks for asking.
answered by: Krutika